OpenStudy (anonymous):

all critical points of 3x(8-x)^3

sam (.sam.):

Critical points means the derivative is 0, You can start by differentiating that then set it equals to zero

OpenStudy (anonymous):

I don't know what to do with the cubed part

sam (.sam.):

Product and chain rule here, --------------------------- Product rule \[\frac{d}{dx}(uv)=udv+vdu\] --------------------------- \[y'=(3x)(3)(8-x)^2(-1)+(8-x)^3(3)\] From chain rule you get (-1) for that derivation. \[y'=-12 (x-8)^2 (x-2)\]

OpenStudy (anonymous):

Thank you!

sam (.sam.):

After that set y'=0 will give you the critical points -12 (x-8)^2 (x-2)=0

OpenStudy (anonymous):

so x= 8 and x=2

sam (.sam.):

Yes